Find the eigenspace corresponding to each eigenvalue of A A 1 3 4 2
Answer #1

Given matrix is As [u ] 2 we find eigenvalues of A consider TA-TI1=0 1-2 3 4 2-d - 10 (1-1) (2-1)-12 =0 -d-21 +22-12=0 12-31② for d=-2 consider (A-II) 2,50 (A +21)2,-0 [:] [:] [:] By :) 1 Ra-URI ] 0 0 >> Hi tha = 0 Take up=t => 1,=-t =-[:] : eigensp